OpenStudy (anonymous):

what is 5/18 + (-5/6)?

OpenStudy (anonymous):

Find a common denominator.

OpenStudy (anonymous):

18

OpenStudy (anonymous):

So then make -5/6 so it has a denominator of 18

OpenStudy (anonymous):

15/18?

OpenStudy (anonymous):

or - 15/18

OpenStudy (anonymous):

Right, -15/18. So now we have \[\large \sf \frac{5}{18}+(-\frac{15}{18})\] which is the same as \[\large \sf \frac{5}{18}-\frac{15}{18}\]

OpenStudy (anonymous):

welp that's probably why I got it incorrect

OpenStudy (anonymous):

Do you understand how to do the rest?

OpenStudy (anonymous):

well tbh now that i'm trying to solve it, no

OpenStudy (anonymous):

\[\large \sf \frac{5}{18}-\frac{15}{18}=\frac{5-15}{18}\]

OpenStudy (anonymous):

10/18?

OpenStudy (anonymous):

Close, but that would be negative. Also, you can simplify it now.

OpenStudy (anonymous):

So what would you get when you simplify \[\large \sf -\frac{10}{18}\]?

OpenStudy (anonymous):

-5/9 right?

OpenStudy (anonymous):

Correct, and that would be your final answer (:

OpenStudy (anonymous):

ok, thank you so much!

OpenStudy (anonymous):

No problem (: